High Voltage set-up in M0’ 2 channels each driving 50 crystals (100 APDs) 1 Crate hosting 1 board (final system). 1 blue cable from Crate to Patch-Panel (could connect up to 9 channels) 2 cables from PP to FFB. Several small cables to distribute the HV to the VFE boards. Connection to the Temperature Interlock from Belgrade University (ramps down HV and then turns off LV)
High Voltage set-up in SM0 34 channels 1 Crate hosting 4 boards 4 blue cables from Crate to Patch-Panel 34 cables from PP to FFB Several small cables to distribute the HV to the VFE boards. Connection to the Temperature Interlock from Belgrade University (ramps down HV and then turns off LV) will switch off the full SuperModule
HV Control and Monitoring PC OPC Server Labview Program
HV Control and Monitoring Program developed by Peter Cogan (Ireland)
HV Control and Monitoring Some problems of comunication between the PC and the OPC server were observed due to the PC being too old or to the software (?) now under investigation… A new PC has been bought dedicated to HV monitoring and control for next year
HV Control and Monitoring The software will be upgraded in 2 phases: 1)Labview program upgrade from 2 channels to 1SM 2)If manpower is sufficient we will develop a test- program with PVSSII
HV stability during 2002 test-beam HV Monitored by the crate is of course very stable 2 mV Monitoring problems not HV problems
¨HV trips¨ in 2002 test-beam During the first week of september a few HV trips occurred. No error status was appearing from the crate. After some investigation we thought it could be noise on the external Temperature-interlock cable (40 metres). After disconnection of the cable no problem was observed for 3 days. So reconnection of the cable with a T+50 Ohm. No problem after 6/9 until end of november Very prompt and kind intervention of CAEN experts to help understand the problem and follow the evolution.
¨HV trips¨ in 2002 test-beam 1.8/8 rack cooling problem 2.14/8 3.16/8 4.28/8 (2) 5.30/8 6.1/9 7.2/9 8.6/9 power cut
Setting time of the APD current When HV is OFF->ON it takes approx. 10 minutes to stabilize the current P3=1/ (h -1 ) Setting time = few minutes Dedicated runs to understand stabilization time of the APD laser signal taken 25/11 will be analized soon
Current during the pion run in-beam current measurements
Conclusions 1.Very good performance of the HV system during the 2002 data-taking 2.In 2003 we will be ready for SM0 run
Your consent to our cookies if you continue to use this website.